### v2

# Detection and Tracking of MAVs Using a Rosette Scanning Pattern LiDAR

Abstract
The use of commercial Micro Aerial Vehicles (MAVs) has surged in the past decade, offering societal benefits but also raising risks such as airspace violations and privacy concerns. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority. In this study, we tackle this challenge, by using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor. The presented method utilizes a particle filter with a velocity component for the detection and tracking of the drone, which offers added re-detection capability. A Pan-Tilt platform is utilized to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense. The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments. Our approach achieved accuracy on par with the state-of-the-art indoor method while increasing the maximum detection range by approximately 80% beyond the state-of-the-art outdoor method.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I INTRODUCTION
Micro Air Vehicles (MAVs) have contributed to the development of many fields, such as precision agriculture [1], modern production [2], transport industry [3] and mining industry [4]. In addition to the above, the utilization of drones (in this paper the phrases drone and MAV are interchangeable) for hobby purposes is also developing explosively, because they have become affordable and easy to handle [5], increasing the risk of inappropriate use. Hobby drones can be used for unauthorized recording, violating people’s right to undisturbed privacy [6, 7]. Furthermore, they can also threaten protected buildings (e.g. airports, factories, power plants) [8]. Due to the increased security risks, the development of autonomous drone detection and tracking systems has become a priority area of research.
In this study, we tackle the challenge of drone detection and tracking using non-repetitive rosette scanning pattern LiDARs, particularly focusing on increasing the detection distance by leveraging the characteristics of the sensor.
We use the rosette scanning pattern Livox AVIA LiDAR sensor, that uses multiple, aligned laser beams. The beams move in a non-repetitive rosette-like pattern inside the sensor, that covers the entire Field of View (FoV) by using Risley prisms [9]. This type of sensor design results in faster measurements, and cheaper devices with fewer moving components compared to the more common rotating models. An important feature of this type of LiDAR is that it captures more measurement data from a target object in the center of the sensor’s FoV, as the laser beams pass through the center each time they travel along a rotating 8-shaped pattern.
Several sensors can be used for drone detection, such as radar [10, 11], LiDAR [12, 13], camera [14, 15, 16], radio-frequency [17, 18] and acoustic sensors [19, 20]. The detection efficiency can be further improved by using sensor fusion [21, 22]. However, each sensor modality should also give good results when used independently [23].
The state-of-the-art drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24]. Dogru et al.[13] achieve a detection distance of $70\;m$ outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/. This detection distance can only be extended to $100\;m$ by adding a retroreflector to the drone. Catalano et al.[24] use a LiVOX Horizon [^5] https://www.livoxtech.com/ to achieve a few $cm$ accuracy using a method based on Kalman Filter which has no re-detection capability in case of lost tracking. Their method was only tested indoors so there is no information on the maximum detection distance.
Our proposed algorithm uses a particle filter with a velocity component to achieve fast detection and tracking with added re-detection capability. Leveraging the point distribution of the LiVOX AVIA LiDAR [^6] https://www.livoxtech.com/avia, the proposed method keeps the MAV in the middle of the sensor’s FoV, where the measurements are most dense, using a pan-tilt platform. With this approach we achieved accuracy on par with the state-of-the-art indoor method using a similar sensor [24], while increasing the maximum detection range approximately $80\%$ beyond the state-of-the-art outdoor method [13] by detecting drones up to $120m$ away without attached retroreflector. In addition, the number of points reflected from the drone at the same distance increased several folds over the full range.
To focus on this challenge, we made two assumptions: (i) the background of the detection area is assumed to be static, and (ii) we disregard the classification problem and assume that all moving objects in the area are drones.
To summarize, this paper enhances the state-of-the-art in LiDAR-based drone tracking with the following contributions:
  • A novel MAV tracking approach for LiDAR utilizing a particle filter with a velocity component is proposed. The updating of particle states of the filter is straightforward using the LiDAR point clouds and the proposed solution offers added re-detection capabilities.
  • Our method uses a Pan-Tilt platform to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense.
  • The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments [^7] Video demonstration of experiments at https://youtu.be/VdePm3on8vo.
The remainder of the paper is organized as follows: related work is presented in section II, the proposed algorithm in section III, the experiments in sections IV and V, and finally the conclusion in section VI.
II RELATED WORK
Tracking MAVs is challenging due to their small size and rapid movement. Traditionally, vision-based approaches have dominated this field. Methods leveraging neural networks, specifically YOLO [25], are widely used for detecting and classifying objects in images and videos. These methods are highly effective in controlled environments with good visibility, as shown in works like [14, 15, 16, 26]. Vision-based techniques leverage large datasets and complex neural network architectures to improve accuracy, but they often struggle in outdoor scenarios where lighting and occlusions affect visibility. Furthermore, the maximum detection distance is not only limited by the resolution of the camera but also the input resolution of the neural network used. In a recent work [27], the maximum detection range of such a system is 40 m.
In response to these limitations, there is a growing interest in non-camera-based solutions, particularly LiDAR, which provides accurate 3D spatial information. LiDAR offers a distinct advantage in detecting and tracking MAVs in low-visibility conditions, where cameras are less effective.
Model-based object detection approaches for LiDAR, such as [28], PointNet [29], PointNet++ [30], VoxelNet [31], and PointPillars [32] are used to detect larger objects like cars or humans, thus making them insufficient for the detection and tracking of drones.
Unlike model-based object detection, model-free methods detect dynamic objects by comparing consecutive scans, making them more versatile. For instance, [33] tracks moving objects by identifying motion patterns between scans, while [34] employs a Kalman filter combined with Iterative Closest Point (ICP) alignment to track objects in dense 3D LiDAR data. Additionally, [35] introduces a method that clusters point clouds using a region growth algorithm and filters objects based on geometric features such as height range and maximum diagonal width. These methods utilize classical rotating LiDARs and also track larger objects like cars or humans.
In [36] a method is proposed for finding MAVs in a point cloud generated by a rosette scanning LiDAR on a moving robot. This approach is further improved by adding an Kalman Filter based tracking method and extensive evaluation in [24]. In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about $30\;m$. However, this method does not allow for regaining tracking once lost. Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor, thereby maximizing the number of returned points. Additionally, our approach incorporates a robust tracking method capable of regaining tracking.
[…]
LiDAR sensor performance in outdoor environments is highly influenced by weather conditions [37], [38]. According to [39] the detection range of LiDAR sensors is limited to $25\;m$ in foggy conditions when the meteorological visibility is under $40\;m$. Based on this a significant reduction in the detection distance and the number of detected points can be expected for outdoor tests under foggy weather conditions. While our experiments confirmed this trend, our method could still track the drone to a little more than $50\;m$.
III THE PROPOSED ALGORITHM
The flowchart of our method is presented in Fig. 1, where the main steps of the algorithm are shown. Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section III-A). This is used to filter the static background from the point cloud to detect the target object.
Figure 1: Overview of the proposed method.
Next, as part of the object tracking, the particle filter[40] performs the prediction step. If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section III-B). The three steps of the particle filter (predict, update, and resample) are explained in detail in Section III-C.
Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section III-D). As a result, the measured point density reflected from the drone is maximized.
III-A Building the Background Model
In our application, the background is considered static throughout the drone tracking, hence it is built once at the beginning of the algorithm. Consequently, the proposed method considers everything that is not recorded in the background a drone. The turret is set to scanning mode and performs movements to scan a pre-defined area. Because of the LiDAR’s non-repetitive way of scanning, the point density of the whole scene will increase the longer the scanning time is.
The background model is represented with OctoMap[41], which is an efficient probabilistic 3D mapping framework using octrees. The octree tree structure results in fast access times when querying individual voxels.
Based on the position and orientation of the LiDAR, the incoming point cloud is transformed into the static coordinate system of the background model. The points are filtered based on position thresholds; the ground plane, near points, and far away points are removed. After filtering, the point cloud is inserted into the OctoMap of the background model. As the last step of building the background model, after inserting all the collected point clouds, inflation is performed to account for measurement noise (e.g. vibrations of the turret base) as well as small changes in background objects (such as leaves moving on the trees).
III-B Processing New Measurements
The point measurements are received in the Cartesian frame of the LiDAR and transformed into the World frame based on the turret’s pan and tilt angles. Then incoming point cloud is filtered as during the background building: the ground plane, near points, and far away points are removed. And finally, the background model is subtracted. There are two additional filters to remove the reflected points that are not close enough to each other and cannot belong to the target object. The filters are the RadiusOutlierRemoval and StatisticalOutlierRemoval algorithms from the Point Cloud Library (PCL) [42].
III-C Object Tracking Using Particle Filter
The drone’s position is estimated using a particle filter, which is a recursive Bayesian state estimator that uses discrete particles to approximate the posterior distribution of the estimated state. It accommodates both linear and non-linear process and measurement models, allowing for flexible, arbitrarily shaped probability distributions. This capability is especially useful for MAVs, which exhibit a wide range of motion patterns with abrupt changes in direction and speed across all three dimensions. Additionally, LiDAR measurements enable straightforward updates of the particle states.
The particle filter operates in three main steps: prediction, update, and resampling. The prediction step uses the current state and transition model to predict the next state. In the update step, incoming filtered point cloud data refines the state estimate via Bayes’ theorem. Finally, resampling aligns the particles with the posterior distribution, addressing the issue of particle degeneracy.
The particle filter is implemented following Algorithm 2. in [43]. Each particle represents a possible state of the drone, where the state is the 3D location and 3D velocity of the drone ($s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$).
III-C1 Predict
First, the position of each particle is predicted based on its previous position, velocity, and added process noise. The prediction noise $p_{\{x,y,z\},k}$ is sampled from a 00 mean Gaussian distribution with $\sigma_{pred}$ standard deviation and $T$ denotes the period of the particle filter.
$p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pred}^{2})$ (1)
$\begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+p_{z,k}\end{bmatrix}$ (2)
The velocities are computed only if the drone is tracked. The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pred}$. The $\sigma_{particles}$ value increases with $\sigma_{pred}$ plus the additional effect of the velocity in every iteration when the drone is lost. The drone is considered tracked if it is under $\sigma_{threshold}=2\cdot\sigma_{pred}$. In this case, the velocities are predicted, or else they are set to 00.
$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}(x_{k}-x_{k-1})/T\\ (y_{k}-y_{k-1})/T\\ (z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
III-C2 Update
The weight of each particle is updated by the product of the previous weight and the Gaussian distribution with $\sigma_{m}$ standard deviation and 00 mean evaluated at $d_{i}$, where $\sigma_{m}$ represents the measurement noise and $d_{i}$ is the distance from the $i$th particle to the closest LiDAR measurement point. This product is normalized, so the sum of $w_{i}$ will add up to $1$, where $n$ is the number of particles.
$w_{i,k}=\frac{w_{i,k-1}\cdot\mathcal{N}(d_{i};0,\sigma_{m}^{2})}{\sum_{i}^{n}w_{i}}$ (4)
III-C3 Resample
After every update step, the particles are resampled using their weights as the probability of being drawn for the next iteration.
III-C4 Particle initialization and reset
The particles are initialized randomly in a cylinder, whose parameters (3D location, radius, and height) are set based on the detection area. The particles can be reset to a new cylinder during runtime, thus giving the operator the capability to reset the tracking to another object. During tracking $\sigma_{particles}<\sigma_{threshold}$ which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$ which is almost certainly a collision of the objects.
III-D Turret control
Turret control is done in two ways: in initialization or tracking mode. During initialization mode, the pan-tilt turret collects data points for a specified time in a predefined area to build the background model. In tracking mode, the control commands are calculated based on the current position of the turret and the estimated position of the drone. The tracking mode aims to keep the drone in the middle of the FoV to maximize the density of reflected points from the drone. In case there are multiple drones in the FoV, the algorithm follows the one closest to the location of the particles, which can be reset by the operator.
III-E Implementation details
The Robot Operating System (ROS) was used to integrate and manage communication between program components, offering a reliable framework for modular nodes. Three nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues, a pan-tilt turret wrapper node to translate MAV position estimates into turret commands, and a MAV detection and tracking node that holds the core algorithm, including the particle filter.
IV INDOOR EXPERIMENTS and EVALUATION
First, the experimental results are presented and examined, starting with the initial detection of the MAV following the construction of the background model. Next, the core of the algorithm, the tracking process, is evaluated in terms of accuracy. Finally, the re-detection capability after tracking loss is assessed.
The experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of $3.5\;m$. Further hardware includes MAVs (Crazyflie 2.0 by Bitcraze), the turret (WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics), and the rosette scanning pattern LiDAR (Avia LiDAR sensor by Livox) with an integration time of $100\;ms$ ($10\;Hz$).
In all our experiments the particle filter was run at $15\;Hz$ frequency which assured the processing of all LiDAR measurements. The processing ran on an 8th-generation desktop Intel i7 processor with less than 30% utilization.
IV-A Detecting the MAV
Figure 2: Initial detection error and $\sigma_{particles}$ of the particle filter. The particle filter locks onto the drone almost instantaneously after receiving the first measurement at around 2 seconds.
Fig. 2 shows the error as the distance of the estimated state to the true state in meters of a representative detection. The average of the standard deviation of the particles in each dimension is also plotted (denoted as $\sigma_{particles}$).
When there is no incoming measurement only the predict step of the particle filter is performed. In all of our experiments $\sigma_{pred}=0.1\;m$ and $n=500$, which turned out to be a good compromise between being able to track fast-moving drones and providing accurate measurements.
While the drone is out of the FoV of the LiDAR sensor, the particles remain idle, only influenced by the added noise in the predict step which results in a slowly inflating particle cloud with a slowly rising $\sigma_{particles}$. In this case, the MAV is resting on the floor and, because of its small size, is hidden inside the background model. As the drone starts hovering and measurements of it arrive at around 2 s the particles gather and allow for the first good estimation. Parallel with this the $\sigma_{particles}$ of the estimation drops sharply. This means the drone is tracked and the particle filter also predicts velocity.
IV-B Tracking the MAV
Figure 3: Different drone testing flight patterns with different colors (vertical, horizontal, fast, and ’lost and found’). The outlined circles show the starting position of each trajectory while the arrows denote the moving direction of the drone. The dashed part of the arrow means that the MAV view is obstructed from the LiDAR’s point of view by an obstacle. The size of each trajectory is in the legend of the figure.
For calculating the accuracy, the MAV was flown following certain flight patterns (described in Fig. 3) in a repeated manner. The ground truth data was obtained by the motion capture system within the MIMO arena by recording the pose of the LiVOX sensor and the position of the drone and aligning the LiVOX measurements with the MIMO arena using the recorded pose. The patterns in Fig. 3 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around $3\;m$. Each circle represents a stationary wait time of $2\;s$ and the one with a black outline represents the first position of the respective pattern. The trajectory represented by the arrows in-between was flown in $3\;s$ in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about $0.6\;m/s$, while it was flown in $2.25\;s$ in case of the fast pattern which resulted in a maximum speed of around $1.2\;m/s$. In each experiment, the drone flies to the first point and then it flies the pattern 3 times in the case of the vertical and horizontal patterns and 4 times in the case of the others.
In Fig. 4, the error of the particle filter is plotted with the ground truth speed calculated from the MIMO measurements. It is visible that the tracking is stable over time with an average position error of under $5\;cm$ in the case of the horizontal and vertical measurements while it is around $7\;cm$ for the fast pattern.
Figure 4: Position error of the particle filter plotted together with ground truth speed from the MIMO arena for the vertical (top), horizontal (middle), and fast (bottom) patterns. The mean ($\mu_{error}$) and standard deviation (±$\pm\sigma_{error}$) of the particle filter error are also plotted.
It can be seen on all the plots of Fig. 4, that the error of the particle filter is dependent on the speed. When the drone is at a waypoint, not moving, the error of the filter is around $3$ to $5\;cm$, while at the peak speeds of $1.2\;m/s$ of the fast pattern (not counting the one outlier), the error is around $20\;cm$. This dynamic difference of about $15\;cm$ in the error can mostly be explained by the different delays in the pipeline. The internal processing time of the LiDAR firmware is $2\;ms$ based on the data-sheet, the LiDAR scan integration time is $100\;ms$, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics. The background subtraction and the application of filters took $10.9\;ms$ with a $\sigma=16.8\;ms$ standard deviation on the fast trajectory and the predict, update and resample steps took a combined red1.$1.11\;ms$ with a $\sigma=1.12\;ms$. The whole delay is about $120\;ms$, in which the drone travels roughly $14.4\;cm$. From the elapsed time, the computation takes only around $20\;ms$ resulting in convenient real-time operation.
The particle filter’s inherent functionality causes the particles to consistently lag slightly behind the measurements. Additionally, the relative motion between the drone and the LiDAR sensor naturally introduces smearing in the point cloud, accounting for the remaining dynamic error.
Overall, when the drone is stationary, the error remains minimal and consistent. However, during movement, a dynamic error arises that increases linearly with speed. This affects the tracking accuracy but does not compromise its robustness.
For comparison, in [24] the RMSE is reported as the metric of accuracy on different trajectories. From their paper, the trajectory T1 is most similar to our horizontal pattern and the trajectory T3 is most similar to our vertical case. We consider them similar because the trajectories are mostly in the horizontal and vertical planes respectively and mostly contain straight lines and turns close to right angles. Their reported best RMSE for T1 and T3 are $0.0773\;m$ and $0.0421\;m$ while the RMSE for our horizontal and vertical patterns are $0.0570\;m$ and $0.0531\;m$ respectively. While the two methods work very similarly in terms of accuracy, our particle filter approach has the added benefit of being able to regain the tracking after it was lost.
IV-C Losing and regaining Tracking
An advantage of utilizing a particle filter is that the detection and tracking properties are part of the same routine, which leads to good performance in terms of regaining tracking after the target is lost.
(a) Stable before loss
(b) Right after loss
(c) Lost for some time
(d) Stable after regain
Figure 5: Different stages of the particle cloud during tracking loss and regain. The particle cloud can be seen with yellow in front of the OctoMap of the background model. The number of particles was 500 in this experiment.
Figure 6: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event. The tracking was lost at around 4 seconds and was regained at 8 seconds.
Fig. 5 shows a typical loss and regain event of the MAV tracking from (a) to (d) while Fig. 6 shows the corresponding error and $\sigma_{particles}$ characteristics. This is one traversal of the drone through the 3 waypoints of the ’lost and found’ pattern. First, the normal MAV tracking is depicted in Fig 5(a). The particles are grouped around the drone with $\sigma_{particles}<0.2\;m$. When the drone starts moving, the error increases slightly but the tracking is still stable. At around 4.5 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $0.2\;m$ meaning the tracking is lost and the particle cloud starts inflating (Fig. 5(b)). Note on the plot, that at first the $\sigma_{particles}$ jumps sharply when the velocity is still estimated and then settles to a much gentler slope when only the prediction noise is added to the particle positions. The particle cloud inflates further to almost $\sigma_{particles}=0.8\;m$ which is depicted in Fig. 5(c) while the error stays constant. This is because the drone hovers behind the obstacle and the estimated pose is the mean of the particle positions which is around the position where the tracking was lost. Then the drone flies out from behind the cover and the particle update and resample steps of the filter condense the cloud instantly which is well depicted on the sharp drop of $\sigma_{particles}$ below $0.2\;m$ in Fig. 6. The tracking is stable again in Fig. 5(d) and the error also decreases to the static error at about 9 seconds when the drone stops moving.
Overall the regaining of tracking after loss is practically instantaneous allowing robust tracking even in scenarios where the trajectory is partially occluded.
V OUTDOOR FIELD TESTS and COMPARISON
After the successful indoor tests, we validated the developed method in an outdoor environment as well. We performed the measurements with the presented method under two different weather conditions (on a sunny and a foggy day). The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in outdoor conditions, on a commercial drone. Finally, we compared the performance of our method to the previous state-of-the-art [13].
V-A Outdoor test
For outdoor tests, the DJI Phantom drone was used on approximately flat terrain. DJI Phantom 4 Pro V2.0 measures $251\;mm\times 398.78\;mm\times 172.72\;mm$ and weighs $1375\;g$ with a battery and propellers attached. An important parameter for detection is the frame size (the largest distance between the motors), which is $350\;mm$ for this type of drone. The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods. For a video of the outdoor experiment, kindly check the video attachment here: https://youtu.be/VdePm3on8vo.
Figure 7: Weather on the test days. Sunny experiment with clouds on the left and the foggy experiment with severely limited viewing conditions on the right.
A sunny day was chosen for the first and a foggy day for the second measurement day when visibility was also severely limited (see Fig.7).
V-B Results
During both tests, the drone was moving under such conditions that both horizontal and vertical distances could be analyzed. In both cases, the controlled pan-tilt turret followed the drone correctly, thus the drone tracking was successful.
(a) Sunny
(b) Foggy
Figure 8: Trajectory and speed of the drone on a sunny and foggy day. The turret was at the origin of the coordinate system.
When evaluating drone detection, an important value is the maximum detection distance. The drone’s trajectory was successfully tracked up to a distance of $128\;m$ in the sunny test environment (Fig. 8(a)), and up to a distance of $58\;m$ in the foggy test environment (Fig. 8(b)). As explained earlier in subsection IV-A, we consider the drone tracked when $\sigma_{particles}<0.2\;m$. We consider the tracking lost after 10 subsequent missing measurements, but it is important to note that we had some occasional detections at distances of up to $145\;m$ on the sunny day.
In both scenarios, we calculated maximum speeds of a little more than $6\;m/s$ from the tracking positions and timestamps.
V-C Comparison
We compare our outdoor test results with another study in which the drone detection efficiency of the VLP-16 LiDAR was tested. In the work of Dogru et al.[13], the size of the drones is larger than the one used for our measurement, which affects the number of reflected points. Furthermore in their paper, there are measurements in which a retroreflector was used to increase the number of points reflected to the sensor.
TABLE I: Comparison of maximum detection distances
Measurement, method Weather With retroreflector Distance($m$)
Our test I. Sunny No 128
Our test II. Foggy No 58
Dogru et al.[13] Sunny Yes 100
Dogru et al.[13] Sunny No 70
Our maximum distance of detection is about $1.5$ to $2$ times greater than their result as in TABLE I.
Another important aspect is the number of detected points, which greatly impacts the robustness of tracking. In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current state-of-the-art (see Fig. 9).
Figure 9: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than [13].
In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[39].
VI CONCLUSION
This article presents a novel tracking method utilizing a rosette scanning pattern LiDAR mounted on a Pan-Tilt platform. The proposed algorithm has been rigorously tested in both indoor and outdoor environments, demonstrating accuracy on par with the current state-of-the-art indoor method [24] and considerably surpassing the leading outdoor method [13] in terms of detection range and point density. Notably, our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects.

### v3

@@
 A pan-tilt platform is utilized to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense.
-The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments.
-Our approach achieved accuracy on par with the state-of-the-art indoor method while increasing the maximum detection range by approximately 80% beyond the state-of-the-art outdoor method.
-This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
-I INTRODUCTION
+The system’s tracking capabilities (both in coverage and distance), as well as its accuracy are validated and compared to State Of The Art (SOTA) models, demonstrating improved performance, particularly in terms of coverage and maximum tracking distance.
+Our approach achieved accuracy on par with the SOTA indoor method while increasing the maximum detection range by approximately $85\;\%$ beyond the SOTA outdoor method to 130m130\;m.
+Additionally, our method yields at least a twofold increase in track coverage and returned point counts.
+Index Terms:
+Drones, Tracking, Trajectory tracking, Particle filters, Sensors
+This work has been published in IEEE ACCES https://doi.org/10.1109/ACCESS.2025.3596857
 Micro Air Vehicles (MAVs) have contributed to the development of many fields, such as precision agriculture [1], modern production [2], transport industry [3] and mining industry [4].
@@
 However, each sensor modality should also give good results when used independently [23].
-The state-of-the-art drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24].
-Dogru et al.[13] achieve a detection distance of $70\;m$ outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/.
-This detection distance can only be extended to $100\;m$ by adding a retroreflector to the drone.
+The State Of The Art (SOTA) drone tracking algorithms using LiDARs either use a classical rotating LiDAR [13] or place a rosette scanning pattern sensor on a stationary mount [24].
+Dogru et al.[13] achieve a detection distance of 70m70\;m outside using a Velodyne VLP-16444https://velodynelidar.com/products/puck/.
+This detection distance can only be extended to 100m100\;m by adding a retroreflector to the drone.
 Catalano et al.[24] use a LIVOX Horizon [^5] https://www.livoxtech.com/ to achieve a few $cm$ accuracy using a method based on Kalman Filter which has no re-detection capability in case of lost tracking.
 Their method was only tested indoors so there is no information on the maximum detection distance.
-Our proposed algorithm uses a particle filter with a velocity component to achieve fast detection and tracking with added re-detection capability.
+Our proposed algorithm uses a Particle Filter (PF) with a velocity component to achieve fast detection and tracking with added re-detection capability.
 Leveraging the point distribution of the LIVOX AVIA LiDAR [^6] https://www.livoxtech.com/avia, the proposed method keeps the MAV in the middle of the sensor’s FoV, where the measurements are most dense, using a pan-tilt platform.
-With this approach we achieved accuracy on par with the state-of-the-art indoor method using a similar sensor [24], while increasing the maximum detection range approximately $80\%$ beyond the state-of-the-art outdoor method [13] by detecting drones up to $120m$ away without attached retroreflector.
-In addition, the number of points reflected from the drone at the same distance increased several folds over the full range.
+With this approach we achieved accuracy on par with the SOTA indoor method using a similar sensor [24], while increasing the maximum detection range approximately $85\%$ beyond the SOTA outdoor method [13] by detecting drones up to 130m130\;m away without attached retroreflector.
+Additionally, our method yields significantly higher track coverage, meaning the drone is tracked for a greater percentage of its trajectory.
+Furthermore, the number of points reflected from the drone at the same distance increased several folds over the full range.
+We provide detailed experiments and comparisons to support these claims and we also analyze how much each component of our algorithm contributes to the observed improvement in tracking distance.
 To focus on this challenge, we made two assumptions: (i) the background of the detection area is assumed to be static, and (ii) we disregard the classification problem and assume that all moving objects in the area are drones.
-To summarize, this paper enhances the state-of-the-art in LiDAR-based drone tracking with the following contributions:
+To summarize, this paper enhances the SOTA in LiDAR-based drone tracking with the following contributions:
   • A novel MAV tracking approach for LiDAR utilizing a particle filter with a velocity component is proposed.
@@
   • Our method uses a pan-tilt platform to take advantage of the specific characteristics of the rosette scanning pattern LiDAR by keeping the tracked object in the center where the measurement is most dense.
-  • The detection capabilities and accuracy of the system are validated through indoor experiments, while the maximum detection distance is shown in our outdoor experiments [^7] Video demonstration of experiments at https://youtu.be/VdePm3on8vo.
-The remainder of the paper is organized as follows: related work is presented in section II, the proposed algorithm in section III, the experiments in sections IV and V, and finally the conclusion in section VI.
-II RELATED WORK
+  • The system’s tracking capabilities (both in coverage and distance), as well as its accuracy are validated and compared to SOTA models [13, 24], demonstrating improved performance, particularly in terms of coverage and maximum tracking distance.
+A video is shared of the experiments.
+[^7] Video demonstration of experiments at https://youtu.be/ghfjqAnDuag
+  • We release the source code of the proposed algorithm and the datasets used in the experiments. [^8] Source code at https://github.com/gsanya/LIVOX_MAV_track
+The remainder of the paper is organized as follows: related work is presented in Section I, the proposed algorithm in Section II, the experiments in Section III, and finally the conclusion in Section IV.
+I RELATED WORK
 Tracking MAVs is challenging due to their small size and rapid movement.
@@
 In a recent work [27], the maximum detection range of such a system is 40 m.
+Other than detection, recent works have also focused on visual tracking.
+The rapid development of deep learning-based visual tracking methods is explored in [28].
+Learned motion models have been shown to outperform traditional Kalman Filter-based approaches, particularly in handling non-linear motion [29].
+However, these methods typically rely on dense visual input and are not well-suited for sparse point clouds with a limited number of measurement points.
+Moreover, these works generally target generic object tracking rather than specifically addressing the challenges of tracking small, fast-moving drones.
 In response to these limitations, there is a growing interest in non-camera-based solutions, particularly LiDAR, which provides accurate 3D spatial information.
@@
 In [38] a method is proposed for finding MAVs in a point cloud generated by a rosette scanning LiDAR on a moving robot.
-This approach is further improved by adding an Kalman Filter based tracking method and extensive evaluation in [24].
-In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about $30\;m$.
+This approach is further improved by adding a Kalman Filter based tracking method and extensive evaluation in [24].
+In these papers, the integration time of the LiDAR is dynamically adjusted to adapt to the speed and distance of the target’s movement enabling the tracking of MAVs indoors up to about 30m30\;m.
 However, this method does not allow for regaining tracking once lost.
-Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor, thereby maximizing the number of returned points.
+Conversely, our aim is to align the rosette scanning pattern LiDAR to ensure continuous centering of the target drone within the sensor FoV, thereby maximizing the number of returned points.
 Additionally, our approach incorporates a robust tracking method capable of regaining tracking.
-[…]
+In the case of a fixed LiDAR configuration, the usability range in [12] is less than 10m10\;m for the VLP16 sensor and 3550m35-50\;m for the HDL-64 sensor.
+In [13], it was demonstrated that few measurements are sufficient to track the targets and the maximum detection range was significantly increased with their method.
+Their approach achieved a maximum detection distance of 70m70\;m with sparse point clouds, and by using retroreflectors, this was extended to 100m100\;m.
+The applied retroreflectors did not increase the size of the drone but increased the number of reflected points from the drone.
+Our method greatly outperforms this baseline solution in similar weather conditions with a maximum tracking distance of more than 130m130\;m without retroreflector.
 LiDAR sensor performance in outdoor environments is highly influenced by weather conditions [39], [40].
-According to [39] the detection range of LiDAR sensors is limited to $25\;m$ in foggy conditions when the meteorological visibility is under $40\;m$.
+According to [41] the detection range of LiDAR sensors is limited to 25m25\;m in foggy conditions when the meteorological visibility is under 40m40\;m.
 Based on this a significant reduction in the detection distance and the number of detected points can be expected for outdoor tests under foggy weather conditions.
-While our experiments confirmed this trend, our method could still track the drone to a little more than $50\;m$.
-III THE PROPOSED ALGORITHM
-The flowchart of our method is presented in Fig.
-1, where the main steps of the algorithm are shown.
-Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section III-A).
+While our experiments confirmed this trend, our method could still track the drone to a little more than 50m50\;m.
+These findings reflect incidental observations rather than a systematic study of weather effects.
+II THE PROPOSED ALGORITHM
+The flowchart of our method is presented in FIGURE 1, where the main steps of the algorithm are shown.
+Note on the figure, when the system is installed at a new location, the first step is to build a background model (see Section II-A).
 This is used to filter the static background from the point cloud to detect the target object.
@@
 Next, as part of the object tracking, the particle filter[42] performs the prediction step.
-If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section III-B).
-The three steps of the particle filter (predict, update, and resample) are explained in detail in Section III-C.
-Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section III-D).
+If there is a new incoming measurement point cloud, the background model is subtracted to reduce static noise (details are given in Section II-B).
+The three steps of the particle filter (predict, update, and resample) are explained in detail in Section II-C.
+Once detected, the drone is kept in the center of the LiDAR’s FoV by orienting the pan-tilt turret towards the position computed by the particle filter (see Section II-D).
 As a result, the measured point density reflected from the drone is maximized.
-III-A Building the Background Model
+II-A Building the Background Model
 In our application, the background is considered static throughout the drone tracking, hence it is built once at the beginning of the algorithm.
-Consequently, the proposed method considers everything that is not recorded in the background a drone.
+Consequently, dynamic objects such as cars, moving foliage, and pedestrians can lead to false positives.
+To address this, we apply simple mitigation strategies that have proven effective in practice.
+While we also explored a transformer-based background segmentation approach in [43], it is not used in this work and falls outside the scope of this paper.
 The turret is set to scanning mode and performs movements to scan a pre-defined area.
@@
 Based on the position and orientation of the LiDAR, the incoming point cloud is transformed into the static coordinate system of the background model.
-The points are filtered based on position thresholds; the ground plane, near points, and far away points are removed.
+The points are filtered based on an operator set bounding box, which is set to remove very far away points, the ground, and ground adjacent points, such as pedestrians and cars.
 After filtering, the point cloud is inserted into the OctoMap of the background model.
-As the last step of building the background model, after inserting all the collected point clouds, inflation is performed to account for measurement noise (e.g.
-vibrations of the turret base) as well as small changes in background objects (such as leaves moving on the trees).
-III-B Processing New Measurements
+As the last step of building the background model, after inserting all the collected point clouds, inflation of the background model is performed to account for measurement noise (e.g.
+vibrations of the turret base) as well as small changes in background objects (such as moving foliage).
+The primary objective of the background removal step is to eliminate the majority of points from the LiDAR measurements that do not correspond to the target.
+Although some false positives may persist, the tracking algorithm has sufficient robustness to handle these residual points.
+II-B Processing New Measurements
 The point measurements are received in the Cartesian frame of the LiDAR and transformed into the World frame based on the turret’s pan and tilt angles.
-Then incoming point cloud is filtered as during the background building: the ground plane, near points, and far away points are removed.
+The incoming point cloud is filtered during the background building: the ground plane, near points, and far away points are removed based on the same bounding box used at background model building.
 And finally, the background model is subtracted.
@@
 The filters are the RadiusOutlierRemoval and StatisticalOutlierRemoval algorithms from the Point Cloud Library (PCL) [45].
-III-C Object Tracking Using Particle Filter
+II-C Object Tracking Using Particle Filter
 The drone’s position is estimated using a particle filter, which is a recursive Bayesian state estimator that uses discrete particles to approximate the posterior distribution of the estimated state.
@@
 in [46].
-Each particle represents a possible state of the drone, where the state is the 3D location and 3D velocity of the drone ($s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$).
-III-C1 Predict
-First, the position of each particle is predicted based on its previous position, velocity, and added process noise.
-The prediction noise $p_{\{x,y,z\},k}$ is sampled from a 00 mean Gaussian distribution with $\sigma_{pred}$ standard deviation and $T$ denotes the period of the particle filter.
-$p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pred}^{2})$ (1)
-$\begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+p_{z,k}\end{bmatrix}$ (2)
-The velocities are computed only if the drone is tracked.
-The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pred}$.
-The $\sigma_{particles}$ value increases with $\sigma_{pred}$ plus the additional effect of the velocity in every iteration when the drone is lost.
-The drone is considered tracked if it is under $\sigma_{threshold}=2\cdot\sigma_{pred}$.
-In this case, the velocities are predicted, or else they are set to 00.
-$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}(x_{k}-x_{k-1})/T\\ (y_{k}-y_{k-1})/T\\ (z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
-III-C2 Update
-The weight of each particle is updated by the product of the previous weight and the Gaussian distribution with $\sigma_{m}$ standard deviation and 00 mean evaluated at $d_{i}$, where $\sigma_{m}$ represents the measurement noise and $d_{i}$ is the distance from the $i$th particle to the closest LiDAR measurement point.
-This product is normalized, so the sum of $w_{i}$ will add up to $1$, where $n$ is the number of particles.
-$w_{i,k}=\frac{w_{i,k-1}\cdot\mathcal{N}(d_{i};0,\sigma_{m}^{2})}{\sum_{i}^{n}w_{i}}$ (4)
-III-C3 Resample
+Each particle represents a possible state of the MAV.
+Three different state representations and transition models were implemented and tested, namely: a delta position based estimator, one using a constant velocity motion model and one using a constant acceleration motion model.
+The state vector $s$ varies depending on the chosen model.
+For the delta position and constant velocity models $s=[x,y,z,v_{x},v_{y},v_{z}]^{T}$ while for the constant acceleration model $s=[x,y,z,v_{x},v_{y},v_{z},a_{x},a_{y},a_{z}]^{T}$.
+The position estimation of the drone is calculated by averaging the particle positions.
+In the upcoming sections we present the equations governing the predict, update, and resample steps of the particle filter for each of the three models.
+II-C1 Predict
+First, the position of each particle is predicted based on its previous position, velocity, acceleration, and added process noise as in (1).
+The position prediction noise $p_{\{x,y,z\},k}$ is sampled from a 0 mean Gaussian distribution with $\sigma_{pos}$ standard deviation, $T$ denotes the period of the particle filter, and $a_{\{x,y,z\},k-1}=0$ in case of the delta position and constant velocity models.
+$\begin{gathered}p_{\{x,y,z\},k}\sim\mathcal{N}(0,\sigma_{pos}^{2})\\ \begin{bmatrix}x_{k}\\ y_{k}\\ z_{k}\end{bmatrix}=\begin{bmatrix}x_{k-1}+v_{x,k-1}\cdot T+0.5\cdot a_{x,k-1}\cdot T^{2}+p_{x,k}\\ y_{k-1}+v_{y,k-1}\cdot T+0.5\cdot a_{y,k-1}\cdot T^{2}+p_{y,k}\\ z_{k-1}+v_{z,k-1}\cdot T+0.5\cdot a_{z,k-1}\cdot T^{2}+p_{z,k}\end{bmatrix}\end{gathered}$ (1)
+The velocities are computed in two different ways.
+In case of the delta position model, the term is dependent on the status of the tracking.
+The tracking can be considered stable when the standard deviation of the particle positions ($\sigma_{particles}$) is around the parameter $\sigma_{pos}$.
+The $\sigma_{particles}$ value increases with $\sigma_{pos}$ plus the additional effect of the velocity in every iteration when the drone is lost.
+The MAV is considered to be successfully tracked if $\sigma_{particles}$ remains below a threshold defined as $\sigma_{threshold}=1.5\cdot\max(\sigma_{pos},\sigma_{m})$, where $\sigma_{m}$ is the measurement uncertainty.
+This condition ensures that when the measurement is less reliable than the position estimate, the resulting estimate reflects the greater uncertainty.
+We scale the velocity with the scaling factor $\rho$ that is inversely proportional to that uncertainty.
+$\rho=\begin{cases}1,&\text{if }\sigma_{particles}<\sigma_{threshold}\\ \frac{\sigma_{pos}}{\sigma_{particles}},&\text{otherwise}\end{cases}$ (2)
+$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}\rho\cdot(x_{k}-x_{k-1})/T\\ \rho\cdot(y_{k}-y_{k-1})/T\\ \rho\cdot(z_{k}-z_{k-1})/T\end{bmatrix}$ (3)
+In case of the constant velocity and constant acceleration models, the velocity is computed as in (4-5), where $a_{\{x,y,z\},k-1}=0$ in case of the constant velocity model.
+$p_{\{v_{x},v_{y},v_{z}\},k}\sim\mathcal{N}(0,\sigma_{vel}^{2})$ (4)
+$\begin{bmatrix}v_{x,k}\\ v_{y,k}\\ v_{z,k}\end{bmatrix}=\begin{bmatrix}v_{x,k-1}+a_{x,k-1}\cdot T+p_{v_{x},k}\\ v_{y,k-1}+a_{y,k-1}\cdot T+p_{v_{y},k}\\ v_{z,k-1}+a_{z,k-1}\cdot T+p_{v_{z},k}\end{bmatrix}$ (5)
+Finally, in case of the constant accelartion model the acceleration is updated as in (6-7).
+$p_{\{a_{x},a_{y},a_{z}\},k}\sim\mathcal{N}(0,\sigma_{acc}^{2})$ (6)
+$\begin{bmatrix}a_{x,k}\\ a_{y,k}\\ a_{z,k}\end{bmatrix}=\begin{bmatrix}a_{x,k-1}+p_{a_{x},k}\\ a_{y,k-1}+p_{a_{y},k}\\ a_{z,k-1}+p_{a_{z},k}\end{bmatrix}$ (7)
+II-C2 Update
+The weight of each particle is updated when a new measurement arrives based on (8).
+The new weight is calculated using a Gaussian distribution with $\sigma_{m}$ standard deviation, which represents the uncertainty of the lidar sensor measurement.
+This distribution is evaluated at $d_{i}$, which is the distance between the $i$th particle and the closest LiDAR measurement point.
+The weights are then normalized (9) so that the sum of $\tilde{w_{i}}$ will add up to 1, where $n$ is the number of particles.
+$w_{i}=\mathcal{N}(d_{i};0,\sigma_{m}^{2})$ (8)
+$\tilde{w_{i}}=\frac{w_{i}}{\sum_{j}^{n}w_{j}}$ (9)
+II-C3 Resample
 After every update step, the particles are resampled using their weights as the probability of being drawn for the next iteration.
-III-C4 Particle initialization and reset
+II-C4 Particle initialization and reset
 The particles are initialized randomly in a cylinder, whose parameters (3D location, radius, and height) are set based on the detection area.
 The particles can be reset to a new cylinder during runtime, thus giving the operator the capability to reset the tracking to another object.
-During tracking $\sigma_{particles}<\sigma_{threshold}$ which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$ which is almost certainly a collision of the objects.
-III-D Turret control
+During tracking $\sigma_{particles}<\sigma_{threshold}$, which means 2 objects are only mistaken if they are closer than $\sigma_{threshold}$, which is almost certainly a collision of the objects.
+II-D Turret control
 Turret control is done in two ways: in initialization or tracking mode.
@@
 In case there are multiple drones in the FoV, the algorithm follows the one closest to the location of the particles, which can be reset by the operator.
-III-E Implementation details
+The pan ($\alpha$) and tilt ($\beta$) goal angles of the turret are calculated based on the current drone position estimate as in (10).
+$\alpha$ $=\tan^{-1}\left(\frac{y_{mav}}{x_{mav}}\right)$ (10)
+$\beta$ $=-\tan^{-1}\left(\frac{z_{mav}}{\sqrt{(x_{mav})^{2}+(y_{mav})^{2}}}\right)$
+II-E Implementation details
 The Robot Operating System (ROS) was used to integrate and manage communication between program components, offering a reliable framework for modular nodes.
-Three nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues, a pan-tilt turret wrapper node to translate MAV position estimates into turret commands, and a MAV detection and tracking node that holds the core algorithm, including the particle filter.
-IV INDOOR EXPERIMENTS and EVALUATION
-First, the experimental results are presented and examined, starting with the initial detection of the MAV following the construction of the background model.
-Next, the core of the algorithm, the tracking process, is evaluated in terms of accuracy.
-Finally, the re-detection capability after tracking loss is assessed.
-The experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of $3.5\;m$.
-Further hardware includes MAVs (Crazyflie 2.0 by Bitcraze), the turret (WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics), and the rosette scanning pattern LiDAR (Avia LiDAR sensor by Livox) with an integration time of $100\;ms$ ($10\;Hz$).
-In all our experiments the particle filter was run at $15\;Hz$ frequency which assured the processing of all LiDAR measurements.
-The processing ran on an 8th-generation desktop Intel i7 processor with less than 30% utilization.
-IV-A Detecting the MAV
+Four nodes were implemented: a LiDAR wrapper node to fix point cloud conversion issues; a pan-tilt turret wrapper node to translate MAV position estimates into turret commands and send them to the turret’s position controller; a background builder node for the Octomap background model; and a MAV detection and tracking node that holds the core algorithm, including the particle filter and other Extended Kalman Filter (EKF) implementations for benchmarking.
+Note that the source code is made available8.
+III Experiments
+In the following subsections, we explain the parameter settings and discuss the experimental results, including a comparison with other methods.
+We first present and validate the equations for setting the parameters of the different motion models of the particle filter in Subsection III-A.
+In Subsection III-B, we quantify the effect of the different parts of our algorithm.
+Then, in Subsection III-C, an extensive comparison is carried out on four recordings, comparing our algorithm with other SOTA methods [13, 24].
+A real-world field test is presented in Subsection III-D.
+The accuracy of the proposed method is investigated in III-E.
+Finally, we show the track regain capability of our algorithm in III-F.
+In all of the experiments, the turret used was the WidowX Dual XM430 Pan & Tilt turret by Trossen Robotics, and the integration time of the LIVOX AVIA sensor was set to 100ms100\;ms.
+III-A Motion model parameter selection
+The particle filter relies on several key parameters: the number of particles $n$ and noise parameters $\sigma_{pos}$, $\sigma_{vel}$, $\sigma_{acc}$, $\sigma_{m}$, which govern how particles evolve and interact with the measurements.
+We justify our parameter choices based on computational constraints, drone dynamics, and experimental validation.
+To ensure real-time operation, we evaluated the effect of particle count $n$ on computational load using the Tracy Profiler [^9] https://github.com/wolfpld/tracy.
+The filter was benchmarked on a single thread of an Intel Core i7 (8th generation) desktop CPU.
+The LiDAR callback can take up a worst-case execution time of 52ms52\;ms (but usually takes much less: 1.45ms1.45\;ms mean with 1.57ms1.57\;ms standard deviation), regardless of the number of particles so each filter step (prediction, update, and resampling combined) had to complete within $48ms$ to maintain a 10Hz10\;Hz filter rate that is needed to process the LiDAR data at 10Hz10\;Hz.
+TABLE I presents the mean, standard deviation, and worst-case execution times of the particle filter across different particle counts and the worst-case full execution time with the LiDAR callback, which has to be under 100ms100\;ms to be able to run at 10Hz10\;Hz.
+TABLE I: Execution times with different particle numbers. Worst-case values are significant outliers, as indicated by the mean and standard deviation.
+n [][-] 10000 9000 8000 7000 6000 5000 4000
+Mean PF $[ms]$ 20.60 17.30 16.03 15.38 12.5 10.85 10.55
+Std PF $[ms]$ 2.52 3.56 1.88 4.54 3.19 3.13 4.81
+Worst-case
+PF $[ms]$
+103.00 88.32 81.38 59.11 50.67 34.62 33.16
+Worst-case
+full $[ms]$
+155.00 140.32 133.38 111.11 102.67 86.62 85.16
+The particles approximate the probability distribution of the model.
+The accuracy of this approximation improves with increasing particle count, asymptotically approaching the true distribution.
+The number was maximized while still enabling real-time operations.
+Based on these results, we set $n=5000$ to ensure real-time performance while maintaining a sufficiently accurate representation of the posterior distribution.
+For context, we report the EKF execution times of the two SOTA methods.
+The mean, standard deviation, and worst-case execution times are (0.12[ms]; 0.11[ms]; 1.73[ms])(0.12\;[ms];\,0.11\;[ms];\,1.73\;[ms]) for [13] and (0.12[ms]; 0.16[ms]; 3.03[ms])(0.12\;[ms];\,0.16\;[ms];\,3.03\;[ms]) for [24].
+While these results indicate noticeably lower execution times, the difference is not significant in the context of the overall pipeline and is justified by the performance improvements outlined in the following Subsections.
+We tested all three motion models for particle prediction.
+Each model predicts particles based on different assumptions and therefore requires different parameter sets.
+Initial tests revealed that while the delta position and constant velocity models performed comparably, the constant acceleration model consistently underperformed even with small $\sigma_{acc}$ values.
+This was due to over-inflation of the particle cloud and poor handling of abrupt directional changes, issues caused by the assumption of continuous acceleration.
+As a result, it was not considered for further evaluation.
+Based on the results of these initial experiments, we derived a unified set of equations for initializing the parameters of all models, grounded in the expected dynamics of the drone.
+This unification promotes consistency and simplifies the tuning process.
+The equations below define $\sigma_{pos}$, $\sigma_{vel}$, $\sigma_{acc}$, and $\sigma_{m}$ in terms of known or measurable drone dynamics, namely expected maximum speed $v_{max}$ and expected maximum acceleration $a_{max}$.
+$\sigma_{pos}$ $=\frac{v_{max}\cdot T}{3}$ (11)
+$\sigma_{vel}$ $=\frac{a_{max}\cdot T}{3}$
+$\sigma_{acc}$ $=0$
+$\sigma_{m}$ $=\frac{s_{diag}}{2}$
+In (11), $T$ is the filter’s time step and $s_{diag}$ is the diagonal distance between the drone’s motors.
+The choice of dividing by 3 ensures that approximately 99.74% of the particles remain within the physically plausible range (assuming a Gaussian distribution).
+The $\sigma_{acc}$ parameter is set to 0 because there is no constantly changing acceleration in the drone’s dynamics.
+Finally, $\sigma_{m}$ was chosen similarly to [24].
+To validate these parameters, we conducted flight experiments with a DJI Phantom 4 ($s_{diag}=0.4\;m$).
+The drone was flown along a challenging trajectory, including high-speed passes perpendicular to the LiDAR’s axis and out-of-range segments to test edge cases that can be seen in FIGURE 2 on the far right.
+Ground Truth (GT) positions were obtained via GPS, streamed in real-time to the processing laptop using UGCS [^10] https://www.ugcs.com/, where the turret was moved based on these GT measurements.
+Despite minor latency, the setup provided reliable ground truth for evaluation.
+Figure 2: Returned point counts and trajectories are shown for the four different evaluation flights.
+The color indicates the number of MAV points in the point cloud from the LIVOX sensor, with red representing zero returns.
+The ’Simple distance’, ’Horizontal slow’, and ’Horizontal fast’ trajectories were flown at constant altitude and are therefore plotted from a top-down perspective.
+The ’Vertical’ trajectory was flown in a line varying the altitude, and is plotted from a side view.
+All trajectories were designed such that, at the maximum distance, no points are returned for extended periods.
+The maximum detection distance is indicated above each figure, with a small black arrow also marking this point along the trajectory.
+The maximum observed velocity and acceleration were approximately 15m/s15\;m/s and $15\;m/s^{2}$, leading to the following parameters based on (11): $\sigma_{pos}=0.5\;m$, $\sigma_{vel}=0.5\;m/s$, $\sigma_{acc}=0\;m/s^{2}$, and $\sigma_{m}=0.2\;m$.
+This speed and acceleration represent the upper limit at which our pilot was still comfortable controlling the drone, and is therefore considered a good approximation of realistic conditions in practical scenarios.
+We then performed a grid search to evaluate the system’s sensitivity to parameter variations:
+[…]
+[…]
+[…]
+[…]
+Each parameter combination was evaluated across 5 trials to account for the non-deterministic nature of the particle filter.
+We used two metrics: (1) maximum tracking distance, defined as the average distance of the furthest correctly identified MAV point cloud (within 4m4\;m of GPS ground truth), and (2) track coverage, the percentage of the flight where the MAV was correctly tracked.
+FIGURE 3 shows a subset of the results for the constant acceleration model for illustration purposes.
+Points are color-coded by the Euclidean distance from the analytical parameter set given by equations (11).
+Results indicate that the filter is robust to small deviations in parameter values, and performance drops only under large perturbations.
+Figure 3: Performance of different parameter combinations in terms of tracking distance and coverage. Colors represent distance from the analytically derived parameter set.
+The narrow performance variance across trials and the existence of a stable parameter zone around the analytical parameter set indicate that the algorithm is stable under parameter perturbations.
+All code, datasets, and evaluation scripts are publicly available at our git repository. This includes raw LiDAR point clouds, GPS ground truth, parameter configuration files, and plotting tools.
+III-B Ablation study
+In this subsection, we evaluate the individual components of our proposed algorithms to assess their contributions to overall performance.
+To isolate the effect of the moving LiDAR, we performed a control experiment where a static, non-moving LIVOX AVIA sensor was positioned next to the moving platform and recorded the same trajectory used in Subsection III-A.
+The number of returned points for both configurations is shown in FIGURE 4, where a significant difference can be observed, particularly in the furthest sections of the trajectory.
+Figure 4: Returned point counts and trajectory for the moving and fixed platforms for the ’Horizontal fast’ trajectory (the left of this figure is the same as the far right of FIGURE 2).
+A notable difference is observed in the furthest sections of the trajectory, where the stationary LiDAR returns almost no points beyond 100m100\;m.
+We compare the proposed delta position and constant velocity motion model PFs with the SOTA method of Dogru et al.
+[13] using an EKF with a constant velocity motion model.
+TABLE II presents the results of our ablation experiments.
+Each experiment was repeated 5 times to account for the non-deterministic nature of the methods.
+The baseline maximum tracking distance of 70m70\;m is the reported result of [13], which was recorded using a circular rotating Velodyne VLP-16 LiDAR that tracked a drone with $s_{diag}=0.45\;m$, slightly larger than our DJI Phantom 4.
+When using the same EKF algorithm within the proposed tracking system, which utilizes the rosette scanning pattern of the LIVOX AVIA sensor, the tracking distance improves by approximately $15\;\%$, which we attribute to the sensor’s superior performance.
+Our particle filters extend this distance by an additional $30\;\%$ in the fixed LiDAR configuration.
+In case of the moving platform, the maximum tracking distances are increased by another $30\;\%$, and are comparable across the EKF and both PFs.
+However, tracking coverage with the PFs is more than five times higher.
+We attribute the additional $30\;\%$ performance gain in maximum tracking distance to our strategy of moving the platform to maintain the drone within the center of the FoV.
+Overall, the best combination of maximum tracking distance and track coverage is achieved by the delta position PF.
+TABLE II: Results of the ablation study.
+Method
+Coverage $\%$]
+mean / std
+Max Dist. $[m]$
+mean / std
+Fix, EKF [13]
+reported with VLP-16
+$N.A.$ $\sim 70/N.A.$
+Fix, EKF [13]
+$6.60/0.03$ $80.3/0.0$
+Fix, PF constant velocity
+$43.91/1.37$ $103.79/0.88$
+Fix, PF delta position
+$51.59/4.47$ $102.83/0.70$
+Moving, EKF [13]
+$12.70/0.42$ $131.70/5.33$
+Moving, PF constant velocity
+$68.42/1.00$ $129.16/0.38$
+Moving, PF delta position
+$71.35/6.07$ $129.68/0.36$
+III-C Tracking performance comparision
+The previously tuned particle filters, utilizing constant velocity and delta position motion models, were benchmarked against the EKF tracking methods presented by Dogru et al.
+[13] and Catalano et al.
+[24].
+These comparisons were conducted within the proposed rosette scanning pattern LiDAR system, where the turret was controlled using GT GPS positions streamed from the drone, ensuring the drone remained centered within the FoV.
+Four different trajectories were considered and recorded on a sunny day.
+These are shown in FIGURE 2, along with the corresponding returned point counts.
+TABLE III summarizes the results of this set of experiments.
+Each test was executed 5 times to account for the non-deterministic nature of the filters.
+We report both the mean and standard deviation for maximum tracking distance and track coverage metrics.
+TABLE III: Maximum detection distance and track coverage results on four outside trajectories.
+Bold font denotes the best result in its column.
+Our PF using the delta position model performs best in terms of coverage in all scenarios, while it matches the SOTA in maximum tracking distance.
+Simple distance Vertical Horizontal slow Horizontal fast
+Method
+Coverage $\%$]
+mean / std
+Max Dist. $[m]$
+mean / std
+Coverage $\%$]
+mean / std
+Max Dist. $[m]$
+mean / std
+Coverage $\%$]
+mean / std
+Max Dist. $[m]$
+mean / std
+Coverage $\%$]
+mean / std
+Max Dist. $[m]$
+mean / std
+EKF [24]
+2.60 / 0.06 14.09 / 0.00 4.41 / 0.21 16.65 / 0.09 2.61 / 0.03 10.88 / 0.32 0.00 / 0.00 0.00 / 0.00
+EKF [24] mod.
+9.83 / 0.21 32.51 / 0.40 20.38 / 0.11 59.42 / 0.55 31.09 / 0.01 71.44 / 0.00 15.20 / 0.00 44.75 / 0.00
+EKF [13]
+37.80 / 1.56 131.23 / 0.00 7.55 / 0.03 51.21 / 0.00 14.08 / 0.03 138.18 / 0.00 12.70 / 0.42 131.71 / 5.33
+PF constant
+velocity
+43.02 / 0.84 131.23 / 0.00 55.84 / 0.51 140.18 / 9.15 73.05 / 0.63 132.02 / 2.28 68.42 / 1.00 129.16 / 0.38
+PF delta position 55.19 / 7.53 131.97 / 2.02 80.35 / 5.25 129.93 / 0.28 74.72 / 3.20 135.57 / 1.61 71.35 / 6.07 129.68 / 0.36
+The EKF in [24] was originally tuned for indoor scenarios with static LiDAR placement, where the drones operate at lower speeds.
+Their method employs a Nearest-Neighbor Search (NNS) algorithm for finding the next possible drone measurement with a fixed radius ($r=s_{diag}/2=0.2\;m$) to associate measurements.
+We benchmarked our algorithm against their original method (EKF [24] in TABLE III) and a modified one, where we increased the NNS search radius (EKF [24] mod.
+in TABLE III).
+It was increased to r=1.5mr=1.5\;m, since this is the expected maximum travel distance based on the drone dynamics in one iteration of the filter.
+The EKF filters were all re-initialized at a specific time, when the LiDAR measurement only contained points from the drone.
+This adjustment was made because their original papers used different approaches for the initial detection, so to ensure a fair comparison, a proper initialization was provided.
+If the drone was successfully tracked for at least five consecutive frames, results are included in the table, otherwise, a value of 0.00 is reported.
+The results indicate that our PF-based algorithms achieve consistently better track coverage than the EKF-based methods.
+The EKF introduced in [24] with the increased NNS radius performs well at close ranges.
+However, once the drone exits the expected radius, the tracker fails and cannot recover.
+[13] partially solves this problem by dynamically setting the NNS radius, which allows their EKF to occasionally reacquire the drone.
+Nevertheless, after that, their EKF prediction is usually inaccurate, and the drone is lost again in a few frames.
+This is reflected in the lower coverage values shown in TABLE III, except for the ’simple distance’ trajectory, where the motion is constant and predictable, enabling successful recovery.
+Our PF-based solutions adapt more effectively to the combination of sensor modality, drone dynamics, and missing measurements, thus having significantly better coverage across all scenarios.
+Due to the unpredictable nature of drone motion, out of the two particle filters, the delta position-based solution performs consistently better in terms of coverage.
+It is somewhat interesting that it also outperforms the model-based one on the ’simple distance’ trajectory, where the speed is almost constant.
+This is likely because the high velocity noise (which is needed to track the drone when it changes direction quickly) results in bigger position errors, ultimately reducing its coverage.
+Overall, we found that our particle filter with delta position motion model matches the SOTA EKF-based solution in terms of maximum tracking distance, while providing significantly better coverage.
+This is particularly important in real-world scenarios, where the turret is moved based on the detection of the filter to keep the drone in the middle of the FoV.
+III-D Field test
+We performed two measurements with the presented delta position PF method under two different weather conditions (on a sunny and a foggy day).
+The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in real-world conditions with turret input from the filter.
+The same DJI Phantom 4 was used for these experiments.
+The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods.
+For a video of the outdoor experiments, kindly check the video attachment here: https://youtu.be/ghfjqAnDuag.
+Figure 5: Weather on the test days. Sunny experiment with clouds on the left and the foggy experiment with severely limited viewing conditions on the right.
+A sunny day was chosen for the first and a foggy day for the second measurement day, when visibility was also severely limited (see FIGURE 5).
+During both tests, the drone was moving under such conditions that both horizontal and vertical distances could be analyzed.
+In both cases, the controlled pan-tilt turret followed the drone correctly, thus the drone tracking was successful.
+(a) Sunny
+(b) Foggy
+Figure 6: Trajectory and speed of the drone on a sunny and foggy day. The turret was at the origin of the coordinate system.
+The drone’s trajectory was successfully tracked up to a distance of 128m128\;m in the sunny test environment (FIGURE 6(a)), and up to a distance of 58m58\;m in the foggy test environment (FIGURE 6(b)).
+As explained earlier in subsection II-C1, we consider the drone tracked when $\sigma_{particles}<0.3\;m$.
+It is important to note that we had some occasional detections at distances of up to 145m145\;m on the sunny day.
+In both scenarios, we calculated maximum speeds of a little more than 6m/s6\;m/s from the tracking positions and timestamps.
+Another important aspect is the number of detected points, which greatly impacts the robustness of tracking.
+In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current SOTA (see FIGURE 7).
+This can be attributed partly to the sensor choice and partly to our strategy of maintaining the drone within the center of the FoV based on our findings in Subsection III-B.
+Figure 7: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than Dogru et al.[13].
+In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[41].
+While a detailed investigation of the performance in different weather conditions is beyond the scope of this work, we include this observation as an incidental but noteworthy outcome of our study.
+III-E Accuracy
+These experiments were conducted within the MIMO arena [26] within the HUN-REN SZTAKI Institute for Computer Science and Control incorporating an OptiTrack motion capture system in a $7\times 8\;m^{2}$ office space with a height of 3.5m3.5\;m, which can provide sub-millimeter accurate, 240 Hz tracking data for bodies with IR retroreflective markers.
+For the target of the experiment a Crazyflie 2.0 MAV by Bitcraze was used, which measures $s_{diag}=12\;cm$.
+The expected maximum velocity and acceleration were 1.2m/s1.2\;m/s and $1.2\;m/s^{2}$ respectively, so the following noise parameters were set for the PF: $\sigma_{pos}=0.04\;m$, $\sigma_{vel}=0.04\;m$, $\sigma_{m}=0.06\;m$.
 Figure 8: Initial detection error and $\sigma_{particles}$ of the particle filter.
 The particle filter locks onto the drone almost instantaneously after receiving the first measurement at around 2 seconds.
-Fig.
-2 shows the error as the distance of the estimated state to the true state in meters of a representative detection.
+$\sigma_{threshold}=\sigma_{m}\cdot 1.5=0.09\;m$ is also plotted.
+FIGURE 8 shows the moment the drone gets tracked by the delta position PF, by plotting the error as the distance of the estimated state to the true state in meters of a representative detection.
 The average of the standard deviation of the particles in each dimension is also plotted (denoted as $\sigma_{particles}$).
-When there is no incoming measurement only the predict step of the particle filter is performed.
-In all of our experiments $\sigma_{pred}=0.1\;m$ and $n=500$, which turned out to be a good compromise between being able to track fast-moving drones and providing accurate measurements.
 While the drone is out of the FoV of the LiDAR sensor, the particles remain idle, only influenced by the added noise in the predict step which results in a slowly inflating particle cloud with a slowly rising $\sigma_{particles}$.
 In this case, the MAV is resting on the floor and, because of its small size, is hidden inside the background model.
-As the drone starts hovering and measurements of it arrive at around 2 s the particles gather and allow for the first good estimation.
-Parallel with this the $\sigma_{particles}$ of the estimation drops sharply.
-This means the drone is tracked and the particle filter also predicts velocity.
-IV-B Tracking the MAV
+As the drone starts hovering and measurements of it arrive at around 2s2\;s the particles gather and allow for the first good estimation.
+Parallel with this, the $\sigma_{particles}$ of the estimation drops sharply.
+This means the drone is tracked.
 Figure 9: Different drone testing flight patterns with different colors (Vertical, Horizontal, Fast, and ’Lost and found’).
@@
 The size of each trajectory is in the legend of the figure.
-For calculating the accuracy, the MAV was flown following certain flight patterns (described in Fig.
-3) in a repeated manner.
+For calculating the accuracy, the MAV was flown following certain flight patterns (described in FIGURE 9) in a repeated manner.
 The ground truth data was obtained by the motion capture system within the MIMO arena by recording the pose of the LIVOX sensor and the position of the drone and aligning the LIVOX measurements with the MIMO arena using the recorded pose.
-The patterns in Fig.
-3 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around $3\;m$.
-Each circle represents a stationary wait time of $2\;s$ and the one with a black outline represents the first position of the respective pattern.
-The trajectory represented by the arrows in-between was flown in $3\;s$ in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about $0.6\;m/s$, while it was flown in $2.25\;s$ in case of the fast pattern which resulted in a maximum speed of around $1.2\;m/s$.
+The patterns in FIGURE 9 were drawn from the point of view of the LiDAR sensor where the middle of each trajectory in the horizontal direction was in front of the sensor with a minimum distance of around 3m3\;m.
+Each circle represents a stationary wait time of 2s2\;s and the one with a black outline represents the first position of the respective pattern.
+The trajectory represented by the arrows in-between was flown in 3s3\;s in case of the vertical, horizontal, and ’lost and found’ pattern which resulted in a maximum speed of about 0.6m/s0.6\;m/s, while it was flown in 2.25s2.25\;s in case of the fast pattern which resulted in a maximum speed of around 1.2m/s1.2\;m/s.
 In each experiment, the drone flies to the first point and then it flies the pattern 3 times in the case of the vertical and horizontal patterns and 4 times in the case of the others.
-In Fig.
-4, the error of the particle filter is plotted with the ground truth speed calculated from the MIMO measurements.
-It is visible that the tracking is stable over time with an average position error of under $5\;cm$ in the case of the horizontal and vertical measurements while it is around $7\;cm$ for the fast pattern.
+TABLE IV: Accuracy comparison for different tracking methods.
+RMSE and 95 % confidence intervals are reported across five runs for each trajectory.
+Our delta position PF consistently achieves the lowest errors, which is particularly apparent in the ‘Fast’ and ‘Lost and found’ scenarios.
+Horizontal Vertical Fast Lost and found
+Method RMSE $[m]$
+CI 95% $[m]$
+lower / upper
+RMSE $[m]$
+CI 95% $[m]$
+lower / upper
+RMSE $[m]$
+CI 95% $[m]$
+lower / upper
+RMSE $[m]$
+CI 95% $[m]$
+lower / upper
+EKF [24]
+0.078 0.076 / 0.081 0.595 0.577 / 0.614 2.604 2.474 / 2.749 1.880 1.810 / 1.955
+EKF [13]
+0.072 0.070 / 0.075 0.071 0.069 / 0.073 0.239 0.227 / 0.252 0.800 0.771 / 0.833
+PF constant
+velocity
+0.080 0.078 / 0.082 0.077 0.075 / 0.080 0.156 0.147 / 0.166 0.643 0.619 / 0.669
+PF delta position 0.054 0.053 / 0.056 0.057 0.055 / 0.059 0.132 0.125 / 0.139 0.292 0.281 / 0.304
+TABLE IV presents a comparison of the accuracy results for the different tracking filters.
+As before, all tests were conducted five times to account for the non-deterministic nature of the filters.
+To ensure a fair comparison, all filters were properly initialized.
+We report the Root Mean Square Error (RMSE) over all five runs, calculated using the Euclidean distance as defined in (13) where $(\hat{x}_{i},\hat{y}_{i},\hat{z}_{i})$ denotes the the filter’s prediction, $(x_{i},y_{i},z_{i})$ denotes the GT position, and $n$ is the number of GT measurements.
+The $95\;\%$ confidence intervals were computed analytically using the Chi-Squared method.
+The ’Lost and found’ trajectory is discussed separately in Subsection III-F.
+$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\sqrt{(\hat{x}_{i}-x_{i})^{2}+(\hat{y}_{i}-y_{i})^{2}+(\hat{z}_{i}-z_{i})^{2}}\right)^{2}}$ (13)
+In these experiments, we also evaluate the proposed method against the EKF implementation from Catalano et al.[24] and the EKF approach proposed by Dogru et al.[13].
+In the case of [24], we used their original NNS search radius, which was tuned for indoor scenarios and did not modify it for this evaluation.
+For the ’Horizontal’ trajectory, the observed results are consistent with those reported in the respective papers.
+However, in the ’Vertical’ and ’Fast’ trajectories, the EKF in [24] performs considerably worse than the other methods.
+This poor performance is likely due to the motion model they employed: a Constant Turn Rate and Velocity (CTRV) model extended with a constant vertical velocity component equal to that of the planar motion.
+This results in degraded estimation along the vertical axis.
+In the ’Fast’ trajectory, tracking is lost almost immediately, which accounts for the high RMSE values there.
+The proposed delta position PF consistently outperforms all other methods across all evaluated scenarios.
+This demonstrates that our approach not only matches the SOTA in terms of maximum tracking distance and surpasses it significantly in terms of track coverage, but is also the most accurate overall.
+In FIGURE 10, the Euclidean error of the delta position particle filter (first of the 5 runs in every scenario) is plotted with the ground truth speed calculated from the MIMO measurements.
+It is visible that the tracking is stable over time with an average position error of around 5cm5\;cm in the case of the horizontal and vertical measurements, while it is around 11cm11\;cm for the fast pattern.
 Figure 10: Position error of the particle filter plotted together with ground truth speed from the MIMO arena for the Vertical (top), Horizontal (middle), and Fast (bottom) patterns.
-The mean ($\mu_{error}$) and standard deviation (±$\pm\sigma_{error}$) of the particle filter error are also plotted.
-It can be seen on all the plots of Fig.
-4, that the error of the particle filter is dependent on the speed.
-When the drone is at a waypoint, not moving, the error of the filter is around $3$ to $5\;cm$, while at the peak speeds of $1.2\;m/s$ of the fast pattern (not counting the one outlier), the error is around $20\;cm$.
-This dynamic difference of about $15\;cm$ in the error can mostly be explained by the different delays in the pipeline.
-The internal processing time of the LiDAR firmware is $2\;ms$ based on the data-sheet, the LiDAR scan integration time is $100\;ms$, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics.
-The background subtraction and the application of filters took $10.9\;ms$ with a $\sigma=16.8\;ms$ standard deviation on the fast trajectory and the predict, update and resample steps took a combined red1.$1.11\;ms$ with a $\sigma=1.12\;ms$.
-The whole delay is about $120\;ms$, in which the drone travels roughly $14.4\;cm$.
-From the elapsed time, the computation takes only around $20\;ms$ resulting in convenient real-time operation.
-The particle filter’s inherent functionality causes the particles to consistently lag slightly behind the measurements.
-Additionally, the relative motion between the drone and the LiDAR sensor naturally introduces smearing in the point cloud, accounting for the remaining dynamic error.
-Overall, when the drone is stationary, the error remains minimal and consistent.
+The mean ($\mu_{error}$) and standard deviation ($\pm\sigma_{error}$) of the particle filter error are also plotted.
+It can be seen on all the plots of FIGURE 10, that the error of the particle filter is dependent on the speed.
+When the drone is at a waypoint, not moving, the error of the filter is around $3$ to 5cm5\;cm, while at the peak speeds of 1.2m/s1.2\;m/s of the fast pattern (not counting the one outlier), the error is around 20cm20\;cm.
+This dynamic difference of about 15cm15\;cm in the error can mostly be explained by the different delays in the pipeline.
+The internal processing time of the LiDAR firmware is 2ms2\;ms based on the data-sheet, the LiDAR scan integration time is 100ms100\;ms, and there is some additional overhead in the formatting, stamping, and publishing of the data to ROS topics.
+The LiDAR callback took 1.45ms1.45\;ms on avarage while the particle filter (predict, update, and resample steps including the background subtraction) took 10.85ms10.85\;ms.
+The whole delay is about 120ms120\;ms, in which the drone travels roughly 14.4cm14.4\;cm, which accounts for most of the dynamic error.
+It is important to note that the size of the Crazyflie drone ($s_{diag}=12\;cm$) is on the same order of magnitude as the observed error.
+The LiDAR only returns points from the side of the drone facing the sensor, and the markers used by the OptiTrack system for tracking are intentionally placed asymmetrically.
+As a result, the true center of the drone does not exactly coincide with the measured GT position.
+Although these factors are difficult to quantify precisely, they all contribute to the overall error.
+Overall, when the drone is stationary, the error remains small and consistent.
 However, during movement, a dynamic error arises that increases linearly with speed.
 This affects the tracking accuracy but does not compromise its robustness.
-For comparison, in [24] the RMSE is reported as the metric of accuracy on different trajectories.
-From their paper, the trajectory T1 is most similar to our horizontal pattern and the trajectory T3 is most similar to our vertical case.
-We consider them similar because the trajectories are mostly in the horizontal and vertical planes respectively and mostly contain straight lines and turns close to right angles.
-Their reported best RMSE for T1 and T3 are $0.0773\;m$ and $0.0421\;m$ while the RMSE for our horizontal and vertical patterns are $0.0570\;m$ and $0.0531\;m$ respectively.
-While the two methods work very similarly in terms of accuracy, our particle filter approach has the added benefit of being able to regain the tracking after it was lost.
-IV-C Losing and regaining Tracking
+III-F Losing and regaining Tracking
 An advantage of utilizing a particle filter is that the detection and tracking properties are part of the same routine, which leads to good performance in terms of regaining tracking after the target is lost.
@@
 The number of particles was 500 in this experiment.
-Figure 6: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event. The tracking was lost at around 4 seconds and was regained at 8 seconds.
-Fig.
-5 shows a typical loss and regain event of the MAV tracking from (a) to (d) while Fig.
-6 shows the corresponding error and $\sigma_{particles}$ characteristics.
+Figure 12: Particle filter error and $\sigma_{particles}$ are plotted together for a typical loss and regain event.
+The error of the EKF of [13] is also shown for comparison.
+The tracking was lost at around 2 seconds and was regained at 6 seconds.
+FIGURE 11 shows a typical loss and regain event of the MAV tracking from (a) to (d) while FIGURE 12 shows the corresponding error and $\sigma_{particles}$ characteristics.
+FIGURE 12 also shows the error of the EKF implementation of [13] which varies the NNS radius based on the uncertainty, giving the method re-detection capability.
 This is one traversal of the drone through the 3 waypoints of the ’Lost and found’ pattern.
-First, the normal MAV tracking is depicted in Fig 5(a).
-The particles are grouped around the drone with $\sigma_{particles}<0.2\;m$.
+First, the normal MAV tracking is depicted in FIGURE 11(a).
+The particles are grouped around the drone with $\sigma_{particles}<\sigma_{threshold}$.
 When the drone starts moving, the error increases slightly but the tracking is still stable.
-At around 4.5 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $0.2\;m$ meaning the tracking is lost and the particle cloud starts inflating (Fig.
-5(b)).
-Note on the plot, that at first the $\sigma_{particles}$ jumps sharply when the velocity is still estimated and then settles to a much gentler slope when only the prediction noise is added to the particle positions.
-The particle cloud inflates further to almost $\sigma_{particles}=0.8\;m$ which is depicted in Fig.
-5(c) while the error stays constant.
-This is because the drone hovers behind the obstacle and the estimated pose is the mean of the particle positions which is around the position where the tracking was lost.
-Then the drone flies out from behind the cover and the particle update and resample steps of the filter condense the cloud instantly which is well depicted on the sharp drop of $\sigma_{particles}$ below $0.2\;m$ in Fig.
-6.
-The tracking is stable again in Fig.
-5(d) and the error also decreases to the static error at about 9 seconds when the drone stops moving.
-Overall the regaining of tracking after loss is practically instantaneous allowing robust tracking even in scenarios where the trajectory is partially occluded.
-V OUTDOOR FIELD TESTS and COMPARISON
-After the successful indoor tests, we validated the developed method in an outdoor environment as well.
-We performed the measurements with the presented method under two different weather conditions (on a sunny and a foggy day).
-The purpose of the validation is to test the track movement and to determine the detection distances and number of detected points in outdoor conditions, on a commercial drone.
-Finally, we compared the performance of our method to the previous state-of-the-art [13].
-V-A Outdoor test
-For outdoor tests, the DJI Phantom drone was used on approximately flat terrain.
-DJI Phantom 4 Pro V2.0 measures $251\;mm\times 398.78\;mm\times 172.72\;mm$ and weighs $1375\;g$ with a battery and propellers attached.
-An important parameter for detection is the frame size (the largest distance between the motors), which is $350\;mm$ for this type of drone.
-The measurements were carried out in the factory state of the drone, we did not improve the detection range by using special paint or other methods.
-For a video of the outdoor experiment, kindly check the video attachment here: https://youtu.be/VdePm3on8vo.
-Figure 7: Weather on the test days. Sunny experiment with clouds on the left and the foggy experiment with severely limited viewing conditions on the right.
-A sunny day was chosen for the first and a foggy day for the second measurement day when visibility was also severely limited (see Fig.7).
-V-B Results
-During both tests, the drone was moving under such conditions that both horizontal and vertical distances could be analyzed.
-In both cases, the controlled pan-tilt turret followed the drone correctly, thus the drone tracking was successful.
-(a) Sunny
-(b) Foggy
-Figure 8: Trajectory and speed of the drone on a sunny and foggy day. The turret was at the origin of the coordinate system.
-When evaluating drone detection, an important value is the maximum detection distance.
-The drone’s trajectory was successfully tracked up to a distance of $128\;m$ in the sunny test environment (Fig.
-8(a)), and up to a distance of $58\;m$ in the foggy test environment (Fig.
-8(b)).
-As explained earlier in subsection IV-A, we consider the drone tracked when $\sigma_{particles}<0.2\;m$.
-We consider the tracking lost after 10 subsequent missing measurements, but it is important to note that we had some occasional detections at distances of up to $145\;m$ on the sunny day.
-In both scenarios, we calculated maximum speeds of a little more than $6\;m/s$ from the tracking positions and timestamps.
-V-C Comparison
-We compare our outdoor test results with another study in which the drone detection efficiency of the VLP-16 LiDAR was tested.
-In the work of Dogru et al.[13], the size of the drones is larger than the one used for our measurement, which affects the number of reflected points.
-Furthermore in their paper, there are measurements in which a retroreflector was used to increase the number of points reflected to the sensor.
-TABLE I: Comparison of maximum detection distances
-Measurement, method Weather With retroreflector Distance($m$)
-Our test I. Sunny No 128
-Our test II. Foggy No 58
-Dogru et al.[13] Sunny Yes 100
-Dogru et al.[13] Sunny No 70
-Our maximum distance of detection is about $1.5$ to $2$ times greater than their result as in TABLE I.
-Another important aspect is the number of detected points, which greatly impacts the robustness of tracking.
-In sunny weather, except for the last part of the detection range, the number of detected points in the result is $3$ to $5$ times higher than in the current state-of-the-art (see Fig.
-9).
-Figure 9: Number of detected points depending on distance. The proposed algorithm returns considerably more points in every distance range than [13].
-In the case of foggy weather, the detection distance was significantly reduced, as expected in advance based on the literature[39].
-VI CONCLUSION
+At around 2 seconds the drone gets behind the obstacle and $\sigma_{particles}$ increases over $\sigma_{threshold}$, meaning the tracking is lost and the particle cloud starts inflating (FIGURE 11(b)).
+Note on the plot that at first the $\sigma_{particles}$ jumps sharply when the $\rho$ ratio is still considerable and then settles to a much gentler slope when only the prediction noise is added to the particle positions.
+The particle cloud inflates further to almost $\sigma_{particles}=0.4\;m$, which is depicted in FIGURE 11(c) while the error stays constant.
+This is because the drone hovers behind the obstacle, and the estimated pose is the mean of the particle positions, which is around the position where the tracking was lost.
+Then the drone flies out from behind the cover, and the particle update and resample steps of the filter condense the cloud instantly, which is well depicted on the sharp drop of $\sigma_{particles}$ below $\sigma_{threshold}$ in FIGURE 12.
+The tracking is stable again in FIGURE 11(d), and the error also decreases to the static error at about 6 seconds when the drone stops moving.
+Overall, the regaining of tracking after loss is practically instantaneous for our PF, while the EKF recovers slightly slower.
+This allows robust tracking even when the trajectory is partially occluded.
+Notably, the PF error is constrained throughout, while the EKF increases constantly until re-detection.
+This difference is clearly illustrated in TABLE IV, where the error of the EKF of [13] is more than twice that of the delta position PF.
+The error reported for [24] is less informative in this context, as their EKF implementation lacks re-detection capability.
+The continuously increasing error of the EKF is not preferred if the turret is moved by the filter prediction, as it may cause the turret to rotate excessively, losing the drone outside the FoV and preventing re-detection altogether.
+For a real-world example where re-detection was demonstrated, kindly check out the attached video7.
+IV CONCLUSION
 This article presents a novel tracking method utilizing a rosette scanning pattern LiDAR mounted on a pan-tilt platform.
-The proposed algorithm has been rigorously tested in both indoor and outdoor environments, demonstrating accuracy on par with the current state-of-the-art indoor method [24] and considerably surpassing the leading outdoor method [13] in terms of detection range and point density.
-Notably, our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects.
+The proposed algorithm has been rigorously tested and compared to SOTA methods in both indoor and outdoor environments, demonstrating accuracy on par with the current SOTA indoor method [24] and surpassing the leading outdoor method [13] by an $85\;\%$ increase in detection range and at least a twofold improvement in both coverage and returned point counts.
+Our method now tracks all dynamic objects, and ongoing work aims to further enhance the system by differentiating between drones and other flying objects.
+Future research directions also include the exploration and integration of deep learning-based tracking techniques, the analysis of environmental effects like fog or rain on our algorithm, and the development of robust multi-object tracking strategies tailored for the pan-tilt configuration.